OpenStudy (mackenzie2013):

Which list contains three equivalent fractions for 2/5 A-4/10, 6/20, 10/25. B- 4/10, 6/15, 8/20 C- 6/15, 8/16, 10/25 D- 6/15, 8/20, 12/25

OpenStudy (shamil98):

Multiply 2/5 by 2/2 what do you get ? ( to find the first fraction )

OpenStudy (mackenzie2013):

I know the first fraction is 4/10 how do I find out the second ?

OpenStudy (shamil98):

Do the same thing to find the second one! 4/10 times 2/2 :)

OpenStudy (mackenzie2013):

Thank you ! :)

OpenStudy (shamil98):

and after you find the second one you do one more multiplication by 2/2 to get the third one.

OpenStudy (skullpatrol):

$$\Huge \frac{a}{b}*\frac{c}{d}=\frac{ac}{bd}$$